a particle is projected from ground in vertical upward direction at t=0 . AT t=0.8 sec ,it reaches h= 14m . it will again come to same height at t= ?

Grade:12

2 Answers

Arun
3 years ago
Dear student
 
h= v₀•t-g•t²/2 , 
v₀=h/t +gt/2 = 14/0.8 +9.8•0.8/2 = 21.42 m/s. 
Time of the motion to the highest point 
t₀=v₀/g =21.42/9.8=2.19 s. 
t1=2•2.19 – 0.8=3.58 s.
Vikas TU
3 years ago